OpenStudy (anonymous):

3/x^2+5x+6 + x-1/x+2 = 7/x+3

OpenStudy (anonymous):

\[\frac{3}{(x+3)(x+2)}+\frac{x-1}{x+2}=\frac{7}{x+3}\]

OpenStudy (anonymous):